OpenStudy (anonymous):

seven more than 3times a number is 31 find the number

OpenStudy (anonymous):

\[3x+7=31\] solve for x by 1) subtracting 7 2) dividing by 3

OpenStudy (anonymous):

3x + 7 = 31. Then you solve for x by subtracting 7 on both side and dividing by 3 on both sides. Looks like this: 3x + 7 = 31 3x +7 -7 = 31 - 7 3x = 24 x = 8